Which ordered pair is a solution to the linear inequality below?
3x-5yis greater than or equal to 10
A. (1,4)
B. (6, -3)
C. (-5, -2)
D. (2, 0)

Answer by jim_thompson5910(35256)

Let's check each answer choice.


Checking choice A. (1,4)

3x-5y >= 10
3*1-5*4 >= 10 ... plug in (x,y) = (1,4)
3-20 >= 10
-17 >= 10 ... false

since -17 >= 10 is false, this makes 3x-5y >= 10 false when (x,y)=(1,4).
(1,4) is NOT a solution.

--------------------------------------------------------------------
Checking choice B. (6,-3)

3x-5y >= 10
3*6-5*(-3) >= 10 ... plug in (x,y) = (6,-3)
18+15 >= 10
33 >= 10 ... true

since 33 >= 10 is true, this makes 3x-5y >= 10 true when (x,y)=(6,-3).
(6,-3) is a solution.
--------------------------------------------------------------------
Checking choice C. (-5, -2)

3x-5y >= 10
3*(-5)-5*(-2) >= 10 ... plug in (x,y) = (-5, -2)
-15+10 >= 10
-5 >= 10 ... false

since -5 >= 10 is false, this makes 3x-5y >= 10 false when (x,y)=(-5, -2).
(-5, -2) is NOT a solution.
--------------------------------------------------------------------
Checking choice D. (2, 0)

3x-5y >= 10
3*(2)-5*(0) >= 10 ... plug in (x,y) = (2, 0)
6-0 >= 10
6 >= 10 ... false

since 6 >= 10 is false, this makes 3x-5y >= 10 false when (x,y)=(2, 0).
(2, 0) is NOT a solution.

==============================================================================================================
Summary:

Only (6,-3) is a solution. So the only answer possible is B. (6, -3)